# Scalable Multi-Objective Reinforcement Learning with Fairness Guarantees using Lorenz Dominance

### v1

@@ 4 Flexible Fairness with $\lambda$-Lorenz Dominance
To give decision-makers fine-grained control over the fairness needs of their specific problem, we introduce a novel criterion, called $\lambda$-Lorenz dominance. $\lambda$-Lorenz dominance operates on the return vectors directly, rather than introducing a weighting over the objectives (such as the Generalized Gini Index (?)). Furthermore, by selecting a single parameter $\lambda\in[0,1]$, $\lambda$-Lorenz dominance allows decision-makers to balance Pareto optimality and Lorenz optimality. We formally define $\lambda$-Lorenz dominance in Definition 3.
@@
$\lambda\sigma({\bm{v}})+(1-\lambda)L({\bm{v}})\succ_{P}\lambda\sigma({\bm{v}}^{\prime})+(1-\lambda)L({\bm{v}}^{\prime}).$ (2)
Intuitively, for $\lambda=1$, it is assumed that the decision-maker cares equally about all objectives and thus may reorder them. This ensures that vectors that were not Pareto-dominated can become dominated. Consider, for example, $(4,2)$ and $(1,3)$. While no vector is Pareto dominated, reordering the objectives in increasing order leads to $(2,4)$ and $(1,3)$ and therefore allows ignoring the second vector. This approach may already reduce the size of the Pareto front, but does not yet achieve the same fairness constraints imposed by the Lorenz front. Setting $\lambda=0$ ensures that the solution set is equal to the Lorenz front.
The $\lambda$-Lorenz front of a given set $D$, denoted $\mathcal{L}\left(D;\lambda\right)$ contains all vectors which are pairwise non $\lambda$-Lorenz dominated. In Theorem 1 we show that, by decreasing $\lambda$, the solution set smoothly converges to the Lorenz front. The detailed proof can be found in the supplementary material.
@@
$\mathcal{L}\left(D\right)\subseteq\mathcal{L}\left(D;\lambda_{1}\right)\subseteq\mathcal{L}\left(D;\lambda_{2}\right)\subseteq\mathcal{F}\left(D\right).$ (3)
@@
${\bm{v}}\succ_{\lambda}{\bm{v}}^{\prime}\implies{\bm{v}}\succ_{L}{\bm{v}}^{\prime}.$ (4)
@@
${\bm{v}}\succ_{\lambda_{2}}{\bm{v}}^{\prime}\implies{\bm{v}}\succ_{\lambda_{1}}{\bm{v}}^{\prime}.$ (5)
@@
Finally, we extend a previous result from ? to show that Pareto dominance implies $\lambda$-Lorenz dominance as well. These components are combined to obtain the desired result. ∎
@@
It is a straightforward corollary that for $\lambda=1,\mathcal{L}\left(D;\lambda\right)\subseteq\mathcal{F}\left(D\right)$ while for $\lambda=0,\mathcal{L}\left(D;\lambda\right)=\mathcal{L}\left(D\right)$. In Figure 2 (right) we illustrate conceptually how the $\lambda$ controls the size of the coverage set to consider.
Figure 2: Lorenz Conditioned Networks (LCNs) is a multi-policy method that offers fair trade-offs between different objectives (left). Reference points enhance training speed by steering training towards desired solution by filtering the Experience Replay buffer (center). $\lambda$-LCN introduces flexibility in fairness preferences, enabling the relaxation of fairness constraints to accommodate more diverse policies (right).
5 Lorenz Conditioned Networks
We introduce Lorenz Conditioned Networks (LCNs), an adaptation of PCNs aimed at learning a $\lambda$-Lorenz front. We then identify a significant drawback in PCN’s management of the ER buffer (Section 5.4) and propose the use of reference points to mitigate ER volatility.
@@ 5 Lorenz Conditioned Networks
LCNs belong to the class of conditioned networks, also referred to as inverse reinforcement learning (IRL), where a policy is trained as a single neural network through supervised learning (?, ?). This policy maps states and desired rewards to probability distributions over actions by using collected experiences—taking actions as labels and states/rewards as inputs. In a multi-objective setting, the network learns multiple policies, each representing a Lorenz-optimal trade-off.
5.1 Network
@@
Just like PCN, LCN uses a single neural network to learn policy $\pi_{\theta}(a_{t}|s_{t},\hat{h_{t}},\ {\hat{\bm{R}}}_{t})$, which maps the current state $s_{t}$, the desired horizon $\hat{h}_{t}$ and the desired return $\hat{\bm{R}}_{t}$ to the next action $a_{t}$. Note that $\hat{\bm{R}}_{t}$ is a vector, with $d$ equal to the number of objectives. The network receives an input tuple $\langle s_{t},\hat{h}_{t},\hat{\bm{R}}_{t}\rangle$ and returns a probability distribution over the next actions. It is trained with supervised learning on samples collected by the agent during exploration. The network updates its parameters using a cross-entropy loss function:
$H=-\frac{1}{N}\sum_{i=1}^{N}\sum_{a\in A}y_{a}^{(i)}\log\pi\left(a_{t}^{(i)}|s_{t}^{(i)},h_{t}^{(i)},\mathbf{R}_{t}^{(i)}\right),$ (6)
@@
where $N$ is the batch size, $y_{a}^{(i)}$ is the $i$-th sample action taken by the agent (ground truth), $y_{a}^{(i)}=1$ if $a_{t}=a$ and 00 otherwise and $\pi(a_{t}^{(i)}|s_{t}^{(i)},h_{t}^{(i)},\mathbf{R}_{t}^{(i)})$ represents the predicted probability of action $a$ for the $i$-th sample, conditioned on its specific state $s_{t}^{(i)}$ , horizon $h_{t}^{(i)}$ , and return $\mathbf{R}_{t}^{(i)}$.
@@
The training process involves sampling collected, non-dominated experiences, and then training the policy with supervised learning to imitate these experiences. Given a sufficient number of good experiences, the agent will learn good policies.
5.2 Collecting experiences
@@
LCN is an off-policy method that learns the policy network $\pi_{\theta}$ by interacting with the environment, collecting experiences, and storing them in an Experience Replay (ER) buffer. These experiences are then used to train the policy via supervised learning (?, ?). Because action selection involves conditioning the network on a specified return, the primary mechanism for collecting higher-quality experiences is to iteratively improve the desired return used as the conditioning input.
@@
To achieve this improvement, a non-dominated return is randomly sampled from the current non-dominated experiences in the ER buffer. This sampled return is then increased by a value drawn from a uniform distribution $U(0,\sigma_{o})$, where $\sigma_{o}$ represents the standard deviation of all non-dominated points in the ER buffer (?). The updated return is subsequently used as the input $\bm{\hat{R_{t}}}$ for the policy network. Through this iterative process—refining the condition, collecting improved experiences, and training the policy network on non-dominated experiences—the network progressively learns to approximate all non-dominated trade-offs, forming a Lorenz coverage set. To ensure that the ER buffer contains experiences that will contribute most to performance improvement, however, the buffer must be filtered to retain only the most useful experiences.
5.3 Filtering experiences
@@
PCN improves the experience replay buffer by filtering out experiences that are far away from the currently approximated Pareto front (non-dominated collected experiences). This is done by calculating the distance of each collected experience to the closest non-Pareto-dominated point in the buffer. In addition, a crowding distance is calculated for each point, measuring its distance to its closest neighbors (?). Points with many neighbors have a high crowding distance and are penalized by doubling their distance metric, ensuring that ER experiences are distributed throughout the objective space (?). With this approach, the set of Pareto-optimal solutions grows exponentially with the number of states and objectives, and maintaining a good ER buffer becomes a great challenge. In applications where fairness is important, much of the Pareto front is undesirable.
Lorenz Conditioned Networks (LCN), on the other hand, seek fairly distributed policies. Thus, the evaluation of each experience $e_{i}$ in the Experience Replay buffer $\mathcal{B}$ is determined by its proximity to the nearest non-Lorenz dominated point $l_{j}\in\mathcal{L}\left(\mathcal{B}\right)\subseteq\mathcal{B}$. In Figure 2 (left) we show an example of this distance calculation. We denote the distance between an experience $e_{i}$ and a reference point $t_{i}$ as $d_{e_{i},t_{i}}=||e_{i}-t_{i}||_{2}$, where $t_{i}=\min||e_{i}-l_{j}||_{2}$ is the nearest non-Lorenz-dominated point, which we refer to as reference point. We formalize the final distance for the evaluation $d_{\text{Lorenz},i}$ as follows:
$d_{\text{Lorenz},i}=\begin{cases}d_{e_{i},t_{i}}&\text{if }d_{cd,i}>\tau_{cd}\\ 2(d_{e_{i},t_{i}}+c)&\text{if }d_{cd,i}\leq\tau_{cd}\end{cases}$ (7)
Where $d_{cd}$ is the crowding distance of $i$ and $\tau_{cd}$ is the crowding distance threshold. A constant penalty $c$ is added to the points below the threshold, whose distance is also doubled. The points in the ER buffer are sorted based on, $d_{\text{Lorenz}}$ and those with the highest get replaced first when a better experience is collected.
5.4 Improving the filtering mechanism with reference points
@@
Consider, for example, vectors: ${\bm{v}}=(8,0),\bm{w}=(3,4)$ and their corresponding Lorenz vectors $L({\bm{v}})=(0,8),L(\bm{w})=(3,7)$. Both ${\bm{v}}$ and $\bm{w}$ are non-Lorenz dominated, and would typically be used as targets for evaluating other experiences. However, ${\bm{v}}$ is not a desirable target due to its unfair distribution of rewards (this is essentially a limitation of Lorenz dominance, when one objective is very large). To address these issues, we adopt the concept of reference points from Multi-Objective Optimization (?, ?, ?), and propose reference points for filtering the experiences. We introduce two reference point mechanisms: a redistribution mechanism and a mean reference point mechanism. Both of these reference points are optimistic, meaning that the agent seeks to minimize the distance to them.
5.4.1 Redistributed Reference Point (LCN-Redist)
@@
$t_{\text{redist}}=\frac{1}{n}\sum_{i=1}^{n}\left(\underset{e\in D}{\mathrm{\operatorname*{arg\,max}}}\sum_{j=1}^{n}e_{j}\right)_{i}$ (8)
Note that $t$ is now the same for all $e\in\mathcal{B}$. Subsequently, we measure the distances of all $e\in\mathcal{B}$ to this reference point and filter out those farthest from it, according to Equation 7 (replace $t_{i}$ with $t_{\text{redist}}$). In Figure 2 (left), we illustrate this transfer mechanism.
5.4.2 Non-Dominated Mean Reference Point (LCN-Mean)
Additionally, we propose an alternative, simpler reference point mechanism: a straightforward averaging of all non-Lorenz dominated vectors in the experience replay (ER) buffer. This approach provides a non-intrusive method for incorporating collected experiences, while simultaneously smoothing out outlier non-dominated points. The reference point, denoted as $t_{\text{mean}}$, is defined as follows: Let $\mathcal{L}\left(B\right)=\{l_{1},l_{2},\ldots,l_{j}\}$ represent the set of non-Lorenz dominated experiences in the ER buffer,
$t_{\text{mean}}=\frac{1}{|\mathcal{L}\left(B\right)|}\sum_{l_{j}\in\mathcal{L}\left(B\right)}l_{j}$ (9)
In Figure 2 (left) we show how this approach defines a reference point.
6 A Large Scale Many-Objective Environment
Existing discrete MORL benchmarks are often small-scale, with small state-action spaces or low-dimensional objective spaces (?, ?). In addition, they do not cover allocation of resources such as public transport, where fairness in the distribution is crucial. We introduce a novel and modular MORL environment, named the Multi-Objective Transport Network Design Problem (MO-TNDP). Built on MO-Gymnasium (?), the MO-TNDP environment simulates public transport design in cities of varying size and morphology, addressing TNDP, an NP-hard optimization problem aiming to generate a transport line that maximizes the satisfied travel demand (?).

### v2

@@ 4. Flexible Fairness with $\lambda$-Lorenz Dominance
To give decision-makers fine-grained control over the fairness needs of their specific problem, we introduce a novel criterion, called $\lambda$-Lorenz dominance. $\lambda$-Lorenz dominance operates directly on the return vectors, without objective weights (such as the Generalized Gini Index (siddique2020learning)). By selecting a single parameter $\lambda\in[0,1]$, $\lambda$-Lorenz dominance allows decision-makers to balance Pareto and Lorenz optimality. We formally define $\lambda$-Lorenz dominance in Definition 1.
@@
(2) $\lambda\sigma({\bm{v}})+(1-\lambda)L({\bm{v}})\succ_{P}\lambda\sigma({\bm{v}}^{\prime})+(1-\lambda)L({\bm{v}}^{\prime}).$
Intuitively, for $\lambda=1$, it is assumed that the decision-maker cares equally about all objectives and thus may reorder them. This relaxation allows some non-Pareto-dominated vectors to become dominated. Consider, for example, $(4,2)$ and $(1,3)$. While no vector is Pareto-dominated, reordering the objectives in increasing order yields $(2,4)$ and $(1,3)$, in which case the second vector is now dominated. This approach may already reduce the size of the Pareto front, but does not yet achieve the same fairness constraints imposed by the Lorenz front. At the other extreme, setting $\lambda=0$ ensures that the solution set is equal to the Lorenz front.
The $\lambda$-Lorenz front of a set $D$, denoted $\mathcal{L}\left(D;\lambda\right)$, contains all vectors that are pairwise non-$\lambda$-Lorenz-dominated. In Theorem 2, we formally show that the $\lambda$-Lorenz fronts form increasing nested sets as $\lambda$ varies from 0 to $1$: lower $\lambda$ values mean increasingly selective coverage sets that interpolate between the Pareto and Lorenz fronts. The full proof is provided in the supplementary material.
@@
(3) $\mathcal{L}\left(D\right)\subseteq\mathcal{L}\left(D;\lambda_{1}\right)\subseteq\mathcal{L}\left(D;\lambda_{2}\right)\subseteq\mathcal{F}\left(D\right).$
@@
(4) ${\bm{v}}\succ_{\lambda}{\bm{v}}^{\prime}\implies{\bm{v}}\succ_{L}{\bm{v}}^{\prime}.$
@@
(5) ${\bm{v}}\succ_{\lambda_{2}}{\bm{v}}^{\prime}\implies{\bm{v}}\succ_{\lambda_{1}}{\bm{v}}^{\prime}.$
@@
Finally, we extend (perny_approximation_2013, Theorem 1) to show that Pareto dominance implies $\lambda$-Lorenz dominance as well. These components are combined to obtain the desired result. ∎
@@
It is a straightforward corollary that for $\lambda=1,\mathcal{L}\left(D;\lambda\right)\subseteq\mathcal{F}\left(D\right)$ while for $\lambda=0,\mathcal{L}\left(D;\lambda\right)=\mathcal{L}\left(D\right)$. In Figure 2 (C) we illustrate conceptually how the $\lambda$ controls the size of the coverage set to consider.
5. Lorenz Conditioned Networks
Lorenz Condition Networks (LCNs) are an adaptation of Pareto Conditioned Networks (PCNs) that aim to learn policies on the $\lambda$-Lorenz front. We use both the abbreviations PCN and PCNs (and likewise LCN and LCNs) interchangeably, depending on whether we refer to the framework as a whole (singular) or to the family of networks (plural). We begin this section by giving an overview of PCNs and identifying their drawbacks.
5.1. Background: Pareto Conditioned Networks (PCN)
A PCN is a supervised learning network designed for Multi-Objective Reinforcement Learning (MORL) (reymond_pareto_2022). It enables a single neural network to learn a diverse set of policies that approximates the Pareto front of optimal trade-offs between objectives. A PCN is a conditioned network, meaning that it is trained to take as input both the environment state and a desired return vector, and output a probability distribution over actions.
Formally, the PCN policy is parameterized as: $\pi_{\theta}(a_{t}\mid s_{t},\hat{h}_{t},\hat{\bm{R}}_{t})$, where $s_{t}$ is the current state, $\hat{h}_{t}$ is the desired horizon, and $\hat{\bm{R}}_{t}$ is a desired return vector over the $d$ objectives. The network is trained using supervised learning on transitions collected during exploration. Each transition includes the state, the action taken, and the return vector obtained from the trajectory, allowing the network to learn by imitating non-dominated behavior conditioned on return goals.
During training, PCN incrementally improves the quality of collected experiences by perturbing non-dominated return vectors in the experience replay (ER) buffer. A new desired return $\hat{\bm{R}}_{t}$ is sampled by adding noise to an existing non-dominated point (proportional to the standard deviation of current returns), which serves as a target for generating new exploratory trajectories. This process helps expand the diversity of collected policies and encourages coverage of the Pareto front.
To filter the ER buffer, PCN introduces a filtering heuristic that favors experiences closer to the Pareto front while preserving diversity. This is achieved by computing the Euclidean distance of each experience to the Pareto front (to prioritize proximity), and a crowding distance (deb_fast_2000) (to encourage spread in objective space), and then applying penalties to overrepresented areas.
Despite its effectiveness in approximating the Pareto front, PCN suffers from two key drawbacks:
  1. (1)
    Lack of fairness control — PCN prioritizes Pareto optimality without considering fairness or equitable reward distribution across objectives, potentially favoring extreme, imbalanced solutions and leading to intractable learning in large state and objective spaces.
  2. (2)
    Experience Replay volatility — As new experiences are collected, many older points are replaced or re-evaluated, causing instability during training.
@@ 5.1. Background: Pareto Conditioned Networks (PCN)
LCNs belong to the same family of reward conditioned networks, also referred to as upside-down reinforcement learning, where a policy is trained as a single neural network through supervised learning (kumar_reward-conditioned_2019; reymond_pareto_2022). An LCN network learns multiple policies, each representing a Lorenz-optimal trade-off.
5.2. LCN Network
@@
Just like PCN, LCN uses a single neural network to learn a policy $\pi_{\theta}(a_{t}|s_{t},\hat{h_{t}},\ {\hat{\bm{R}}}_{t})$, which maps the current state $s_{t}$, the desired horizon $\hat{h}_{t}$ and the desired return $\hat{\bm{R}}_{t}$ to the next action $a_{t}$. Note that $\hat{\bm{R}}_{t}$ is a vector with dimension $d$ equal to the number of objectives. The network receives an input tuple $\langle s_{t},\hat{h}_{t},\hat{\bm{R}}_{t}\rangle$ and returns a probability distribution over the set of potential next actions. It is trained with supervised learning on samples collected by the agent during exploration. The network updates its parameters using a cross-entropy loss function:
(6) $\mathcal{J}(y,\pi)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{a\in A}y_{a}^{(i)}\log\pi\left(a_{t}^{(i)}|s_{t}^{(i)},h_{t}^{(i)},\bm{R}_{t}^{(i)}\right),$
@@
where $N$ is the batch size, $y_{a}^{(i)}$ is the $i$-th sample action taken by the agent (ground truth), $y_{a}^{(i)}=1$ if $a_{t}=a$ and 0 otherwise and $\pi(a_{t}^{(i)}|s_{t}^{(i)},h_{t}^{(i)},\bm{R}_{t}^{(i)})$ represents the predicted probability of action $a$ for the $i$-th sample, conditioned on its specific state $s_{t}^{(i)}$, horizon $h_{t}^{(i)}$, and return $\bm{R}_{t}^{(i)}$.
@@
The training process involves sampling collected, non-dominated experiences and then training the policy with supervised learning to imitate these experiences. Given a sufficient number of good experiences, the agent will learn good policies.
5.3. Collecting Experiences
@@
LCN learns the policy network $\pi_{\theta}$ by collecting experiences and storing them in an ER buffer. These experiences are then used to train the policy via supervised learning (kumar_reward-conditioned_2019; reymond_pareto_2022). Because action selection involves conditioning the network on a specified return, the primary mechanism for collecting higher-quality experiences is to iteratively improve the desired return used as the conditioning input.
@@
To achieve this improvement, similarly to PCN, a non-dominated return is randomly sampled from the current non-dominated experiences in the ER buffer. This sampled return is then increased by a value drawn from a uniform distribution $U(0,\sigma_{o})$, where $\sigma_{o}$ represents the standard deviation of all non-dominated points in the ER buffer (reymond_pareto_2022). The updated return is subsequently used as the input $\bm{\hat{R_{t}}}$ for the policy network. Through this iterative process—refining the condition, collecting improved experiences, and training the policy network on non-dominated experiences—the network progressively learns to approximate all non-dominated trade-offs, forming a Lorenz coverage set. To ensure that the ER buffer contains experiences that will contribute most to performance improvement, however, the buffer must be filtered to retain only the most useful experiences.
5.4. Filtering Experiences
@@
PCN improves the experience replay buffer by filtering out experiences that are far away from the currently approximated Pareto front. This is done by calculating the distance of each collected experience to the closest non-Pareto-dominated point in the buffer. In addition, a crowding distance is calculated for each point, measuring its distance to its closest neighbors (deb_fast_2000). Points with many neighbors have a high crowding distance and are penalized, ensuring that ER experiences are distributed throughout the objective space (reymond_pareto_2022). With this approach, the set of Pareto-optimal solutions can grow exponentially with the number of states and objectives (perny_approximation_2013), and maintaining a good ER buffer becomes a great challenge. This can be intuitively understood by a simple example provided in (perny_approximation_2013): consider a deterministic MOMDP with $N+1$ states, where each non-terminal state allows two actions with different two-objective reward vectors. The terminal state is absorbing, and assume a discount factor $\gamma=1$. There are $2^{N+1}$ possible stationary deterministic policies, and exactly $2^{N}$ of them result in distinct value vectors at the initial state. These vectors take the form $(x,2^{N}-1-x)$ for $x=0,1,\dots,2^{N}-1$, and all lie on the Pareto front since improving one objective necessarily worsens the other. Therefore, the number of Pareto-optimal value vectors grows exponentially with $N$, the number of non-terminal decision points.
This explosion is a consequence of the definition of Pareto dominance, which treats all objectives as equally important and assumes no prior knowledge of user preferences. Any improvement in any objective is valued, resulting in a broad set of solutions. However, in many practical applications, such a solution set may be unnecessary or even undesirable. For instance, when objectives correspond to benefits for distinct individuals or groups (and we know that fairness is preferred), some Pareto-optimal solutions may be undesirable, as they result in highly unequal outcomes.
It is important to note that this issue is not exclusive to Pareto dominance; other dominance relations, such as Lorenz dominance, can theoretically also produce exponentially large solution sets. Empirically, for a fixed number of objectives, approximating a Lorenz-optimal set within an error bound from the Lorenz front can be computed in polynomial time on the size of the state and objective spaces. We provide empirical results supporting this claim in Section 7. This aligns with the intuition that stricter dominance criteria, such as those that prioritize fairness, tend to reduce the number of acceptable solutions by implicitly encoding preference information without relying on explicit weights. This distinction is crucial for experience filtering: when prior knowledge about user preferences (e.g., fairness) is available, employing stricter dominance criteria can produce more computationally manageable ER buffers.
[…]
(7) $d_{\text{Lorenz},i}=\begin{cases}d_{e_{i},t_{i}}&\text{if }d_{cd,i}>\tau_{cd}\\ \rho_{pen}(d_{e_{i},t_{i}}+c)&\text{if }d_{cd,i}\leq\tau_{cd}\end{cases}$
Where $d_{cd}$ is the crowding distance of $i$ and $\tau_{cd}$ is the crowding distance threshold. A constant penalty $c$ is added to the points below the threshold, whose distance is additionally penalized by a penalty multiplier $\rho_{pen}$. For the experiments in this paper, we set $\rho_{pen}=2$ and provide additional sensitivity analysis on Appendix D. The points in the ER buffer are sorted based on $d_{\text{Lorenz}}$, and those with the highest get replaced first when a better experience is collected. The threshold $\tau_{cd}$ separates serves to break high-density regions in the objective space and ensures diversity. Experiences with a crowding distance below this threshold lead to too similar policies and are therefore penalized. While this heuristic introduces a discontinuity at the threshold $\tau_{cd}$, in practice this acts as a strong bias toward maintaining diversity in the buffer, which benefits stable training.
Illustration of Lorenz Conditioned Networks (LCNs) as a multi-policy method for fair trade-offs between objectives. Panel A shows standard LCNs balancing objectives by exploring the trade-off space. Panel B shows how reference points guide learning by filtering the Experience Replay buffer and reducing outlier influence. Panel C shows lambda-LCNs, which introduce flexibility in fairness preferences to allow more diverse policies. Each group corresponds to an objective related to group satisfaction, such as satisfied transportation demand.
Figure 2. Lorenz Conditioned Networks (LCNs) is a multi-policy method that offers fair trade-offs between different objectives. In this illustrative example, each group (Group 1 and Group 2) corresponds to a distinct objective related to group satisfaction (e.g. satisfied transportation demand). (A) Standard LCNs learn policies that balance objectives by exploring the trade-off space between them. (B) Reference points help accelerate training by filtering the Experience Replay buffer and guiding learning toward desirable solutions, while reducing the influence of outliers. (C) $\lambda$-LCN introduces flexibility in the fairness preferences, enabling the relaxation of fairness constraints to accommodate more diverse policies.
5.5. Improving the Filtering Mechanism with Reference Points
@@
Consider, for example, vectors: ${\bm{v}}=(8,0),\bm{w}=(3,4)$ and their corresponding Lorenz vectors $L({\bm{v}})=(0,8),L(\bm{w})=(3,7)$. Both $\bm{v}$} and $\bm{w}$ are non-Lorenz-dominated, and would typically be used as targets for evaluating other experiences. However, $\bm{v}$} is not a desirable target due to its unfair distribution of rewards (this is essentially a limitation of Lorenz dominance, when one objective is very large). To address these issues, we adopt the concept of reference points from Multi-Objective Optimization (deb_moo_refpoints; cheng_ref_vector; felten_decomposition), and propose reference points for filtering the experiences. We introduce two reference point mechanisms: a redistribution mechanism and a mean reference point mechanism. Both of these reference points are optimistic, meaning that the agent seeks to minimize the distance to them.
5.5.1. Redistributed Reference Point (LCN-Redist)
@@
(8) $t_{\text{redist}}=\frac{1}{n}\sum_{i=1}^{n}\left(\underset{e\in D}{\mathrm{\operatorname*{arg\,max}}}\sum_{j=1}^{n}e_{j}\right)_{i}.$
Note that $t$ is now the same for all $e\in\mathcal{B}$. Subsequently, we measure the distances of all $e\in\mathcal{B}$ to this reference point and filter out those farthest from it, according to Equation 7 (replace $t_{i}$ with $t_{\text{redist}}$). In Figure 2 (B), we illustrate this transfer mechanism.
5.5.2. Non-Dominated Mean Reference Point (LCN-Mean)
Additionally, we propose an alternative, simpler reference point mechanism: a straightforward averaging of all non-Lorenz-dominated vectors in the experience replay (ER) buffer. This approach provides a non-intrusive method for incorporating collected experiences, while simultaneously smoothing out outlier non-dominated points. The reference point, denoted as $t_{\text{mean}}$, is defined as:
(9) $t_{\text{mean}}=\frac{1}{|\mathcal{L}\left(B\right)|}\sum_{l_{j}\in\mathcal{L}\left(B\right)}l_{j},$
where $\mathcal{L}\left(B\right)=\{l_{1},l_{2},\ldots,l_{j}\}$ represent the set of non-Lorenz-dominated experiences in the ER buffer. In Figure 2 (B) we show how this approach defines a reference point.
6. A Large Scale Many-Objective Environment
Existing discrete MORL benchmarks are often small-scale, with small state-action spaces or low-dimensional objective spaces (vamplew_empirical_2011; lopez_microscopic_2018). In addition, they do not cover allocation of resources such as public transport, where fairness in the distribution is crucial. We introduce a novel and modular MORL environment, named the Multi-Objective Transport Network Design Problem (MO-TNDP). Built on MO-Gymnasium (Alegre+2022bnaic), the MO-TNDP environment simulates public transport design in cities of varying sizes and morphologies, addressing TNDP, an NP-hard optimization problem aiming to generate a transport line that maximizes the satisfied travel demand (farahani2013review).